OpenStudy (anonymous):

If ƒ(x ) = x 2 + 1 and g(x ) = 3x + 1, find ƒ(2) + g(3). 5/3 12 15 what is the answer

OpenStudy (mathstudent55):

Evaluate the function f at 2 and the function g at 3. Then add the results. \(f(x) = x^2 + 1\) and \(g(x) = 3x + 1\) \(f(2) = 2^2 + 1\) and \(g(3) = 3(3) + 1\) \(f(2) = 4 + 1 = 5\) and \(g(3) = 9 + 1 = 10\) \(f(2) + g(3) = 5 + 10 = 15\)